# Scalable mixed-domain Gaussian process modeling and model reduction for longitudinal data

### v2

@@ Abstract
Gaussian process (GP) models that combine both categorical and continuous input variables have found use in longitudinal data analysis of and computer experiments. However, standard inference for these models has the typical cubic scaling, and common scalable approximation schemes for GPs cannot be applied since the covariance function is non-continuous. In this work, we derive a basis function approximation scheme for mixed-domain covariance functions, which scales linearly with respect to the number of observations and total number of basis functions. The proposed approach is naturally applicable to also Bayesian GP regression with discrete observation models. We demonstrate the scalability of the approach and compare model reduction techniques for additive GP models in a longitudinal data context. We confirm that we can approximate the exact GP model accurately in a fraction of the runtime compared to fitting the corresponding exact model. In addition, we demonstrate a scalable model reduction workflow for obtaining smaller and more interpretable models when dealing with a large number of candidate predictors.
@@ 1 Introduction
Gaussian processes (GPs) offer a flexible nonparametric way of modeling unknown functions. While Gaussian process regression and classification are commonly used in problems where the domain of the unknown function is continuous, recent work has seen use of GP models also in mixed domains, where some of the input variables are categorical or discrete and some are continuous. Applications of mixed-domain GPs are found e.g. in Bayesian optimization (Garrido-Merchán and Hernández-Lobato, 2020), computer experiments (Deng et al., 2017; Roustant et al., 2020; Wang et al., 2021; Zhang and Notz, 2015) and longitudinal data analysis (Cheng et al., 2019; Timonen et al., 2021). For example in biomedical applications, the modeled function often depends on categorical covariates, such as treatment vs. no treatment, and accounting for such time-varying effects is essential. Non-trivial mixed-domain kernels are required in order to obtain an interpretable model with both shared and category-specific effects (Timonen et al., 2021).
@@ 4.3.1 Eigendecomposition
Let us study a kernel $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{V}\subset\mathbb{R}$, where $\mathcal{X}$ is a finite set of $C$ possible values (categories). We can encode these categories numerically as integers $\mathcal{X}=\{1,\ldots,C\}$. Because there are only $C^{2}$ possible input combinations $(v,w)$ for $k$, and therefore $\lvert\mathbb{V}\rvert\leq C^{2}$, we can list them in the $C\times C$ matrix $\bm{C}$ which has elements $\left[\bm{C}\right]_{v,w}=k(v,w)$. If $k$ is symmetric, the symmetric square matrix $\bm{C}$ has the orthogonal eigendecomposition
@@ 6.4 Experiment 4: Model reduction with simulated data
An additional experiment with a larger data set and different number of true variables is described in Sec. S2.
@@ 7 Conclusion
Gaussian processes offer an attractive framework for specifying flexible models using a kernel language. The computational cost of their exact inference however limits possible applications to small data sets, and the complexity of alternative model spaces poses additional limits for the scalability of a useful model building workflow. Our scalable framework opens up a rich class of GP models to be used in large scale applications of various fields of science as the computational complexity is linear with respect to data size. We have presented a scalable approximation scheme for mixed-domain covariance functions, and demonstrated its use in the context of Bayesian GP regression. However, it can also be applied in GP applications where the kernel hyperparameters are optimized using a marginal likelihood criterion. A limitation is that we have assumed that the categorical kernels are symmetric, and continuous kernels are stationary. Non-stationary effects can still be modeled by applying a warping on the input first, and then using a stationary kernel (see for example (Snoek et al., 2014)).
@@
[Acknowledgments] We thank Aki Vehtari and Gleb Tikhonov for useful comments on early versions of this manuscript, and acknowledge the computational resources provided by Aalto Science-IT, Finland.

### v3

@@ Abstract
Gaussian process (GP) models that combine both categorical and continuous input variables have found use in analysis of longitudinal data and computer experiments. However, standard inference for these models has the typical cubic scaling, and common scalable approximation schemes for GPs cannot be applied since the covariance function is non-continuous. In this work, we derive a basis function approximation scheme for mixed-domain covariance functions, which scales linearly with respect to the number of observations and total number of basis functions. The proposed approach is naturally applicable to also Bayesian GP regression with discrete observation models. We demonstrate the scalability of the approach and compare model reduction techniques for additive GP models in a longitudinal data context. We confirm that we can approximate the exact GP model accurately in a fraction of the runtime compared to fitting the corresponding exact model. In addition, we demonstrate a scalable model reduction workflow for obtaining smaller and more interpretable models when dealing with a large number of candidate predictors.
@@ 1 Introduction
Gaussian processes (GPs) offer a flexible nonparametric way of modeling unknown functions. While Gaussian process regression and classification are commonly used in problems where the domain of the unknown function is continuous, recent work has seen use of GP models also in mixed domains, where some of the input variables are categorical or discrete and some are continuous. Applications of mixed-domain GPs are found e.g. in Bayesian optimization (Garrido-Merchán and Hernández-Lobato, 2020), computer experiments (Qian et al., 2008; Deng et al., 2017; Roustant et al., 2020; Wang et al., 2021; Zhang and Notz, 2015) and longitudinal data analysis (Cheng et al., 2019; Timonen et al., 2021). For example in biomedical applications, the modeled function often depends on categorical covariates, such as treatment vs. no treatment, and accounting for such time-varying effects is essential. Non-trivial mixed-domain kernels are required in order to obtain an interpretable model with both shared and category-specific effects (Timonen et al., 2021).
@@ 3 Related research
Computer experiments
In analysis of computer experiments (Sacks et al., 1989), GPs can be used as a surrogate model for the output of a computer code that takes a long time to run. Often, both qualitative and quantitative factors affect the output of a computer code. Whereas the kernels used in this paper are valid by construct, a strategy to find general valid covariance functions for computer experiments involving both continuous and categorical variables was developed by Qian et al. (2008), using semidefinite programming. The output of the computer code can be assumed to be observed exactly, meaning that there is no model for the observation noise (Sec. 2.2). In this case, the presented approximation methodology can still be used via the strategy detailed in Sec. S3.3. with $\sigma^{2}=0$, as no noise can be seen as a special case of the Gaussian observation model with $\sigma^{2}\rightarrow 0$.
@@ 4.3.1 Eigendecomposition
Let us study a kernel $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{V}\subset\mathbb{R}$, where $\mathcal{X}$ is a finite set of $C$ possible values (categories). We can encode these categories numerically as integers $\mathcal{X}=\{1,\ldots,C\}$. Because there are only $C^{2}$ possible input combinations $(v,w)$ for $k$, and therefore $\lvert\mathbb{V}\rvert\leq C^{2}$, we can list them in the $C\times C$ matrix $\bm{C}$ which has elements $\left[\bm{C}\right]_{v,w}=k(v,w)$. The symmetric square matrix $\bm{C}$ has the orthogonal eigendecomposition
@@ 4.5 Mixed kernels for longitudinal data
4.6 Motivation and advice for base kernels
The used base kernels $k_{j,q}$, $\kappa_{j,r}$ should be selected so that they reflect the assumed covariance structure over the dimension of x in question. In this paper we use the exponentiated quadractic kernel since we have observed it to work well enough, and its spectral density has a known closed form expression. The practitioner who wishes to use a different kernel needs to derive or look up its spectral density (Eq. 12). See for example (Rasmussen and Williams, 2006) for the formula for the spectral density of the Matern class of covariance functions.
For categorical variables we have used the zero-sum kernel because it solves a non-identifiability problem of additive population-level and group-level effects, as the group level effects sum to zero when the ZS kernel is used. Furthermore, the eigendecomposition (Eq. 16) of any kernel of the compound symmetry class has a convenient closed form expression (Sec. 4.3.2) and the ZS kernel is a valid compound symmetry kernel by definition. For a different categorical kernel, one may need to restrict its possible parameters so that positive semidefiniteness is ensured, and compute the corresponding eigendecomposition.
@@ 6.4 Experiment 4: Model reduction with simulated data
An additional experiment with a larger data set and different number of true variables is described in Sec. S2.1.
7 Limitations
The approximation method that we use requires that the continuous kernels are stationary. Non-stationary effects can still be modeled by applying a warping on the input first, and then using a stationary kernel (see for example Snoek et al. (2014)). Another limitation is that we use kernels that are composed of sums and products of base kernels. It is possible to construct a mixed-domain kernel that does not fit into this family, as is done in Sec. S2.2, where we study one such kernel. However, we find this example rather artificial, and in reality additive kernels with low-dimensional base kernels have demonstrated good predictive performance (Lu et al., 2022), and have the advantage of increased model interpretability.
In Sec. 6, the largest data size that was analyzed was $N=12775$ observations. In Sec. S2.3, we push the modeling framework further to study the practical scalability limit. We find estimate that with our implementation, it is viable to analyze data sets on the order of $N=10^{5}$ when the total number of basis functions $M$ is on the order of hundreds (see Sec. 4.1). This practical limit arises largely due to using full Bayesian inference with MCMC, as Stan does not scale well to millions of observations.
@@ 8 Conclusion
Gaussian processes offer an attractive framework for specifying flexible models using a kernel language. The computational cost of their exact inference however limits possible applications to small data sets, and the complexity of alternative model spaces poses additional limits for the scalability of a useful model building workflow. Our scalable framework opens up a rich class of GP models to be used in large scale applications of various fields of science as the computational complexity is linear with respect to data size. We have presented a scalable approximation scheme for mixed-domain covariance functions, and demonstrated its use in the context of Bayesian GP regression. However, it can also be applied in GP applications where the kernel hyperparameters are optimized using a marginal likelihood criterion.
@@
[Acknowledgments] We thank Aki Vehtari and Gleb Tikhonov for useful comments on early versions of this manuscript, and acknowledge the computational resources provided by Aalto Science-IT, Finland. We wish to also thank anonymous reviewers for their constructive comments.